OpenStudy (anonymous):

[8.03] Using complete sentences, explain how to completely factor 3x2 − 3x − 60. Help please.! :)

OpenStudy (anonymous):

no

jhonyy9 (jhonyy9):

for the first step you can factorizing out from evra 3 terms 3 ,yes ? so than will get 3(x^2 -x -20) = so than now we need working with x^2 -x -20 so you see that the coeficient of second term is -1 what is b if you consider the standard form of quadratic equation ax^2 +bx +c =0 and -20 = c so for you can factorizing this quadratic you need getting 2 numbers p and q with property that p-q=-1 and p*q=-20 if you check some numbers you will get that for p=-5 and q=4 you will get the necessary results for b and c so -5+4=-1 and -5*4=-20 so hence you can rewriting this quadratic factorized in the form of x^2 -x -20 = (x-5)(x+4) but don't forget it that the initially quadratic was 3x^2 -3x -60 = 3(x^2 -x -20) = 3(x-5)(x+4) so hope so much that you have understood it right,easy good luck bye

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

yw was my pleasure good luck hope so much that is understandably sure easy right